OpenStudy (anonymous):

simplify the complex fraction: 1-(1/4) over 2-(3/5) I don't understand what i'm doing wrong on this one.

jimthompson5910 (jim_thompson5910):

1 - 1/4 = 4/4 - 1/4 = (4-1)/4 = 3/4 So 1 - 1/4 = 3/4 Do the same with 2 - 3/5 and tell me what you get

OpenStudy (anonymous):

but the example in the book is showing me something completely different and the answer says the whole thing is 15/28 an i don't understand how they got the answer. when i tried it like the book shows me i get a different answer

jimthompson5910 (jim_thompson5910):

Let me show you how they got that answer

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

You agree that 1 - 1/4 = 3/4 right?

OpenStudy (anonymous):

yeh

jimthompson5910 (jim_thompson5910):

Let's now simplify 2 - 3/5

jimthompson5910 (jim_thompson5910):

2 - 3/5 10/5 - 3/5 (10-3)/5 7/5 So 2 - 3/5 = 7/5

jimthompson5910 (jim_thompson5910):

With me so far?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

So 1-(1/4) over 2-(3/5) turns into (3/4) over (7/5)

jimthompson5910 (jim_thompson5910):

Now we just need to simplify (3/4) over (7/5)

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

(3/4) over (7/5) (3/4)*(5/7) (3*5)/(4*7) 15/28 So (3/4) over (7/5) = 15/28 Which means that 1-(1/4) over 2-(3/5) fully simplifies to 15/28

OpenStudy (anonymous):

okay so hold on. you keep change flip and then cross multiply?

jimthompson5910 (jim_thompson5910):

yes, to divide two fractions, you flip the second one and you multiply

jimthompson5910 (jim_thompson5910):

you multiply numerators together and denominators together

OpenStudy (anonymous):

oh okay you multiply the numerators and then the denominators. right?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Okay, I think I've got it now! Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

so wait how would you do (1/3)/(3/b)?

jimthompson5910 (jim_thompson5910):

To divide two fractions, flip the second fraction "3/b" to get "b/3" then multiply So (1/3)/(3/b) becomes (1/3)*(b/3)

jimthompson5910 (jim_thompson5910):

then multiply numerators: 1 and b ---> 1*b = b and the denominators: 3 and 3 ---> 3*3 = 9 So (1/3)*(b/3) = b/9 which means (1/3)/(3/b) = b/9

OpenStudy (anonymous):

oh. Thanks.

jimthompson5910 (jim_thompson5910):

sure thing